### v2

@@ 7 Conclusion
When applied to price vanilla European options with cash dividends under the piecewise-lognormal model (by reducing the problem to a vanilla basket option pricing problem), the expansions are found to be the most precise when compared to previously published approximations tailored to the specific case of cash dividends.
@@ Appendix A Derivatives of the Black-76 formula towards the strike
$\frac{\partial}{\partial K}\mathrm{Black}(F,K,v,T)$ $=-\eta B(0,T)\Phi(\eta d_{2})\,,$ (20)
$\frac{\partial^{2}}{\partial K^{2}}\mathrm{Black}(F,K,v,T)$ $=\frac{B(0,T)}{K\sqrt{v}}\phi(d_{2})\,,$ (21)
$\frac{\partial^{3}}{\partial K^{3}}\mathrm{Black}(F,K,v,T)$ $=\frac{B(0,T)}{K^{2}\sqrt{v}}\left(\frac{\phi(d_{2})}{\sqrt{v}}-1\right)\,,$ (22)

### v3

@@ 7 Conclusion
When applied to price vanilla European options with cash dividends under the piecewise-lognormal model (by reducing the problem to a vanilla basket option pricing problem), the expansions are found to be the most precise against previously published approximations tailored to the specific case of cash dividends.
@@ Appendix A Derivatives of the Black-76 formula towards the strike
(20) $\frac{\partial}{\partial K}\mathrm{Black}(F,K,v,T)$ $=-\eta B(0,T)\Phi(\eta d_{2})\,,$
(21) $\frac{\partial^{2}}{\partial K^{2}}\mathrm{Black}(F,K,v,T)$ $=\frac{B(0,T)}{K\sqrt{v}}\phi(d_{2})\,,$
(22) $\frac{\partial^{3}}{\partial K^{3}}\mathrm{Black}(F,K,v,T)$ $=\frac{B(0,T)}{K^{2}\sqrt{v}}\left(\frac{\phi(d_{2})}{\sqrt{v}}-1\right)\,,$